ACME Credit Union has $200,000 in RSAs earning an average yield of 4.25%. ACME has RSLs of $150,000 at a cost of .50%.

ACME has a                            [ Select ]                       ["negative", "positive"]         gap.

What is ACME's expected annual Net Interest Income?                            [ Select ]                       ["7,750", "8,500", "5,625", "750", "7,500"]      

If interest rates increased by 1% for both assets and liabilities, what would be the expected annual Net Interest Income?                            [ Select ]                       ["2,250", "7,500", "5,625", "815", "8,250"]      
Answer #1

1: Positive

GAP=RSA-RSL = 200000-150000=$50000

2: 7750

expected annual Net Interest Income= 4.25%*200000-0.5%*150000=$7750

3: 8250

expected annual Net Interest Income= 5.25%*200000-1.5%*150000=$ 8250